OpenStudy (anonymous):

if both the roots of x^2-(p-4)x+2e^(2log p)-4=0 are negative then what are the possible range of values of p

OpenStudy (anonymous):

x^2-(p-4)x+2P^2-4=0

OpenStudy (anonymous):

if both are -ve so p-4<0 and p<4 also p>0 as it is in ln.

OpenStudy (amistre64):

i was on the cusp of saying something similar :)

OpenStudy (anonymous):

but agin diccrim. will be +ve.

OpenStudy (anonymous):

why should p-4 be less than zero if the roots are negative?

OpenStudy (amistre64):

(r+m)(r+n) amkes the roots both -

OpenStudy (anonymous):

(p-4)^2>4(2p^2-4) 7p^2+8p-32<0

OpenStudy (amistre64):

+m+n = + result for -(p-4)

OpenStudy (amistre64):

-p+4 >/ 0 -p >/ -4 p </ 4

OpenStudy (amistre64):

2e^(2log p)-4 >/ 0 since +m(+n) = +result

OpenStudy (anonymous):

not following you guys..:-(

OpenStudy (anonymous):

what is your prob?????????

OpenStudy (amistre64):

suppose we got: (r-4)(r-6) are the zeros here + or - ?

OpenStudy (anonymous):

dipankar: u said "if both are -ve so p-4<0" why is this so?

OpenStudy (anonymous):

because if ax^2+bx+c=0 has roots m,n then m+n=-b/a.

OpenStudy (anonymous):

here b=-(p-4), a=1 and two roots are-ve so m+n=p-4 is -ve.

OpenStudy (amistre64):

2e^(2log p)-4 >/ 0 2e^(2log p) >/ 4 e^(2log p) >/ 2 ln(2 log(p)) >/ ln(2) ln(2) + ln(log(p)) >/ ln(2) ln(log(p)) >/ 0

OpenStudy (anonymous):

ok...got it now.. so whats nxt?

OpenStudy (amistre64):

ln(ln(0)/ln(10)) >/ 0 ln(ln(p)) - ln(10) >/ 0 i think this is how to do it :)

OpenStudy (anonymous):

p<4 also p>0 is that the range?

OpenStudy (anonymous):

no you also have to satisfy that D>=0

OpenStudy (anonymous):

i.e. 7p^2+8p-32<0

OpenStudy (anonymous):

if this gives some value of p that is less than 4 but>0 that will be your range.

OpenStudy (amistre64):

ln(ln(0)/ln(10)) >/ 0 ln(p)/ ln(10) >/ e^0 = 1 ln(p) >/ ln(10) p >/ 10 ok, what if anything did i do wrong

OpenStudy (amistre64):

ln(p)/ ln(10) >/ e^0 = 1 log(p) >/ 1 ....same result for me that way.....

OpenStudy (anonymous):

what is your problem now?????????

OpenStudy (anonymous):

a<0, b<0 a+b<0, ab>0 therefore p-4<0 => p<4; 2p^2-4>0 =>p^2>2 therefore p>sqrt{2} or, p<-sqrt(2) so, p\[p \in(\sqrt{2},4) [p>0]\] again both roots are -ve. b^2-4ac>=0 (p-4)^2-4(2p^2-4)>=0 =>7p^2+8p-32<=0 therefore \[\sqrt{2}<p<(4/7)(\sqrt{15}-1)\] did i get it right?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

what are the roots of the eqn 7p^2+8p-32<=0

OpenStudy (anonymous):

can you tell me please.

OpenStudy (anonymous):

(4(+-)sqrt(230))/7

OpenStudy (anonymous):

how????????? its wrong.

OpenStudy (anonymous):

(-8(+-)sqrt(960))/7

OpenStudy (anonymous):

yea i factored 2 out

OpenStudy (anonymous):

n yea it should be -ve 4 up there

OpenStudy (anonymous):

hey how did u get ur denominator to be 7?? shouldnt it be 14??

OpenStudy (anonymous):

sorry 14.

OpenStudy (anonymous):

and also it sqrt(920) nt 960..

OpenStudy (anonymous):

why? 32*4*7+64

OpenStudy (anonymous):

yea..ryt..my bad

OpenStudy (anonymous):

wrong!, what is all this noobery

OpenStudy (anonymous):

it says both the roots are negative , not that they are complex

OpenStudy (anonymous):

it says both the roots are negative , not that they are complex

OpenStudy (anonymous):

if both the roots are negative that means that the sum must be negative, and the product must be positive